OpenStudy (anonymous):

When solving negative one over five (x − 25) = 7, what is the correct sequence of operations? Multiply each side by negative one over five , add 25 to each side Multiply each side by 5, subtract 25 from each side Multiply each side by negative one over five , subtract 25 from each side Multiply each side by −5, add 25 to each side

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

we have to multiply both sides by \(-5\), so we can write this: \[\left( { - 5} \right) \cdot \left( { - \frac{1}{5}} \right) \cdot \left( {x - 25} \right) = 7 \cdot \left( { - 5} \right)\]

OpenStudy (anonymous):

so it would be B

OpenStudy (michele_laino):

since: \[\left( { - 5} \right) \cdot \left( { - \frac{1}{5}} \right) = 1\]

OpenStudy (michele_laino):

please note that I have multiplied by \(-5\)

OpenStudy (anonymous):

So D

OpenStudy (michele_laino):

correct!